OpenStudy (anonymous):

help please solve the system using substitution x=3y-4 x+y=16

jimthompson5910 (jim_thompson5910):

x+y=16 3y-4+y=16 ... replace x with 3y-4 4y-4 = 16 y = ???

OpenStudy (anonymous):

would you subtract 4 from 16 to make it 12 or 4 from 4 ?

jimthompson5910 (jim_thompson5910):

you add 4 to both sides actually

jimthompson5910 (jim_thompson5910):

you do that to move the "-4" over so to speak

OpenStudy (anonymous):

oh yeah. so move "-4" over to 16?

jimthompson5910 (jim_thompson5910):